OpenStudy (princeevee):

which expresion can be added to the right side of the equal sign to form an equation? 42 - 5x2 = ? A. (5+3)x 2 B. 2(6+1) C. 5x8-2 D. 4x9+2

OpenStudy (calculusxy):

Find out the value of the expression that they gave you and then find another expression in the choices that has the same value.

OpenStudy (princeevee):

but i cant get it well : /

OpenStudy (calculusxy):

42 - (5 x 2) = ?

OpenStudy (princeevee):

c?

OpenStudy (calculusxy):

This questions looks a bit weird :\ 42 - (5 x 2) = 42 - 10 = 32 A. (5+3) x 2 = 15 x 2 = 30 B. 2(6+1) = 2(7) = 14 C. 5 x 8 - 2 = 40 - 2 = 38 D. 4 x 9 + 2 = 36 + 2 = 38

OpenStudy (princeevee):

so whats the answer?

OpenStudy (calculusxy):

I am not sure !

OpenStudy (princeevee):

could you ask someone else?

OpenStudy (princeevee):

to help?

OpenStudy (calculusxy):

Yeah

OpenStudy (calculusxy):

@e.mccormick Can you please check out this question ?

OpenStudy (e.mccormick):

What it looks like it is asking is what has the same meaning as 42 - 5x^2

OpenStudy (princeevee):

can you see the answer?

OpenStudy (e.mccormick):

Nonthing there seems to the same... hmmm.

OpenStudy (princeevee):

OOOO

OpenStudy (princeevee):

42 is supose to be 24

OpenStudy (calculusxy):

Oh...

OpenStudy (princeevee):

mispell )=)=)B D

OpenStudy (calculusxy):

24 - 10 = 14 So look at the answer choice that has the same answer ! :)

OpenStudy (princeevee):

which is?

OpenStudy (e.mccormick):

So 24 - 5x^2 = ? cause the x2 looks like it is supposed to be x^2.

OpenStudy (princeevee):

no its multiplication there

OpenStudy (calculusxy):

I actually gave the answer to you...

OpenStudy (princeevee):

im just a deer troll dont judge me

OpenStudy (calculusxy):

Look back at the other replies

OpenStudy (e.mccormick):

Aaah, \(\times\) not ex.... umm... what grade is this for?

OpenStudy (princeevee):

9th

OpenStudy (calculusxy):

truthfully, this does not look like a ninth grade problem to me.

OpenStudy (princeevee):

it is